simplify (show your work)
3/5 + 1 and 1/5 - 1 and 2/3

Answer by fractalier(2101)

Okay from
3/5 + 1 1/5 - 1 2/3
let us change everything to the lowest common denominator, which here is 15ths...so we get
3/5 + 1 1/5 - 1 2/3 =
9/15 + 1 3/15 - 1 10/15 =
1 12/15 - 1 10/15 =
2/15